Mzshay438:

Determine the value of . ​ ​Select the two answers from the choices. A 45∘ B 75∘ C 50 D 135∘ E 90∘ E 90∘90^{\circ}90 ∘

T21KSavage:

whats the question?

T21KSavage:

determine the value of what?

RomanSeals:

What choices do u think u can eliminate?

Extrinix:

90 x 90 to the 90th power i think?

Mzshay438:

How do I send the picture

RomanSeals:

now multiply 9 times 9 and leave the zero

Mzshay438:

Mzshay438:

Here's the pic

Mzshay438:

I don't understand it

RomanSeals:

What grade are u in?

Mzshay438:

I'm in the 11th but flunk 10th grade

Mzshay438:

Math

Mzshay438:

Can somebody help

Mzshay438:

Before I go banannas

Mzshay438:

girlygirl:

mmm 90 i think sorry if it wrong

jhonyy9:

so you need find the value of x and the y ? like a first step hope you know that the angles measures of a triangle equal 180 degrees using this you can write x +x +2x = 180 4x = 180 x = 180/4 x = ? and again what you need to know that a strait angle - so a line - has 180 degrees to find the y value just subtract the x value from 180 and will get sure easy the y value hope helped understandably easy @supie

Mzshay438:

Thanks

Mzshay438:

So the 1st question is 45

Mzshay438:

Answer

Mzshay438:

How do you find the y again I'm lost